OpenStudy (1018):

Justin is riding a circular ferris wheel with a radius of 5.50 m. If it completes two revolutions every minute, what is its centripetal acceleration?

ganeshie8 (ganeshie8):

`it completes two revolutions every minute` How much time does it take to complete one revolution ?

OpenStudy (1018):

half a minute?

OpenStudy (1018):

30 s

ganeshie8 (ganeshie8):

which is same as 30 seconds

ganeshie8 (ganeshie8):

\(\omega = \dfrac{2\pi}{T} = \dfrac{2\pi}{30}\) rad/sec

ganeshie8 (ganeshie8):

next use the formula centripetal acceleration = \(\omega^2r\)

OpenStudy (christos):

you are gonna have to make use of 2 formulas for circular motion

OpenStudy (1018):

hey can you help i cant seem to get the answer. :(

OpenStudy (1018):

having a hard time with the rev/min

OpenStudy (christos):

Well you are gonna have to convert rev/min into rev/sec

OpenStudy (christos):

before you do anything else

OpenStudy (christos):

then you have to convert use that to find T (period)

OpenStudy (christos):

T = 1 revolution